Find an equation (s) of the parabola (s) with horizontal or vertical axis, vertex (2,-3) and containing (6,-1).
Answer by lwsshak3(11628)

Find an equation (s) of the parabola (s) with horizontal or vertical axis, vertex (2,-3) and containing (6,-1).
***
Vertex form of equation for:
..
Vertical axis of symmetry: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex
Solving for A:
-1=A(6-2)^2-3
-1=16A-3
16A=2
A=2/16=1/8
Equation:
y=(1/8)(x-2)^2-3
..
Horizontal axis of symmetry: x=A(y-k)^2+h, (h,k)=(x,y) coordinates of the vertex
Solving for A:
6=A(-1-(-3))^2+2
6=A(-1+3))^2+2
6=4A+2
4A=4
A=1
Equation:
x=(y+3)^2+2